OpenStudy (anonymous):

Solve for x: −3(x + 3) = −3(x + 1) − 5

OpenStudy (konradzuse):

ok so lets start distributing first. -3(x+3) = -3x -9

OpenStudy (konradzuse):

On the other side we have -3x-3 -5

OpenStudy (anonymous):

-3x-9=-3x-3-5 -3x-9=-3x-8 -6x=1 x=1/-6

OpenStudy (anonymous):

its not x=1/6

OpenStudy (konradzuse):

Yeah that's because he did -3x -3x to be -6x, but it would be -3x + 3x which should be 0x, so I'm konfused myself lol.

OpenStudy (konradzuse):

Are you sure it's -3 for both?

OpenStudy (konradzuse):

Otherwise there shouldn't be a solution to this....

OpenStudy (anonymous):

if you wrote the problem correctly then the answer i provided is correct however if not then the answer must have no solution.

OpenStudy (konradzuse):

-3x = -3x you would have to add the 3x to both sides, or you could divide -3x/-3 and get x = x x - x = .... 0 = ....

OpenStudy (anonymous):

6

OpenStudy (konradzuse):

@lgbasallote any clue to this mystery?

OpenStudy (anonymous):

No solution

OpenStudy (konradzuse):

so no solution was correct then eh? I win :)